At what distance of separation must two 1.00-microCoulomb charges be positioned in order for the repulsive force between them to be equivalent to the weight (on Earth) of a 1.00-kg mass?
Answer #1

F = 1*9.8 = kq^2/d^2 = 8.98*10^9*10^-12/d^2 = 8.98*10^-3/d^2
d = 3.027 cm